# “Talking Past Each Other”: Issue Ownership and Microtargeting in Swiss Online Political Ads

### v1

@@ 4 RESULTS
4.1 Microtargeting
Figure 2 compares each party’s audience demographics to the “party complementary” profile, which consists of all advertisements not associated with that party. Since RQ1.a does not concern regional differences, ads from all languages are combined for this analysis. Gender disparities are quantified using odds ratios (OR), which measure the increased likelihood that an ad impression will be viewed by a man. An OR of 1 represents perfect gender balance; an OR less than 1 indicates a female bias; and an OR greater than 1 indicates a male bias. For age, the divergence between distributions is measured using the Wasserstein distance (WD), which indicates the mean difference in years for a party versus the complementary. Overall, centrist and right-wing parties are more oriented towards a male audience, particularly the SVP ($36\%$). On the other hand, the GRÜNE and the SP are more female-oriented, with the GRÜNE at $27\%$ and the SP at $13\%$. In terms of age reach, the SVP and, to a lesser extent, Die Mitte tend to attract an older audience, whereas the GRÜNE and the GLP primarily attract people in the 25–34 age group. Furthermore, the GRÜNE is the only party that extensively reaches the 18–24 age group, yet it underrepresents the over-45 age group more than any other party.
Figure 2: Comparison of the demographic reach by age and gender of each party’s ads versus all other ads. Gender is measured using male-to-female odds ratios, and age using Wasserstein distances.
Figure 3: Language-specific age distributions (bootstrapped mean and SD) for each party’s ads compared to its complementary ad set. The Wasserstein Distance measures the divergence between the two.
Figure 3 shows the age distribution of the audience reached by the parties, compared to the party complementaries, broken down by language group. The shaded area represents the standard deviation of the audience reached for each age group. The distribution is build using a bootstrapping procedure similar to that described for Figure 4. Significant differences between parties within the same canton can be observed, as well as regional variations within parties. While the SVP focuses on an audience over 45, especially in Ticino, Die Mitte focuses on an older audience in French-speaking cantons only. The SP, in German-speaking cantons, has a WD of almost 0 (no differences with the complementary party’s audience), while in French-speaking cantons, it targets a much more over-45 audience. In Ticino, however, it does the opposite, underrepresenting the over-35 audience and focusing on the 18-24 age range. The GLP, on the other hand, focuses heavily on the 25-34 age range in French-speaking cantons, while in Ticino, the reach is broader (25-45). With the exception of the GRÜNE and the SVP, parties in the German-speaking cantons show fewer differences compared to their counterparts. In summary, parties differ in their demographic reach (RQ1.a), with evident internal variations emerging across different linguistic regions (RQ1.b).
Figure 4: Male-to-female odds ratios for each party, relative to its complement, by language. Values are bootstrapped means, and 1 means perfect gender balance. An asterisk indicates that the 95% confidence interval does not include 1.0.
4.2 Referenda
@@
Figure 1 shows the seven-day moving average of total political ad views. The gray vertical lines represent referendums, and the inset shows the total number of views per language in the 30 days before the federal elections of October 22, 2023. Clearly, this event caused the most significant increase in views. At first glance, it is instead difficult to answer RQ2.a and determine whether advertising activity increases before the 14 referendum dates. As preliminary step, a Wilcoxon signed-rank test [Wilcoxon] is applied to compare ad views and spending in the 30 days before and after each referendum. For completeness, this analysis is conducted separately for German, French, and Italian ads. The results of all the tests (shown in Table  A5) confirm a significant increase in ad views and spending in the 30 days before a referendum ($p<0.001$ in all cases).
@@
Although the Wilcoxon signed-rank test confirms a difference, it does not account for confounding factors, such as seasonality or other temporal patterns. The true impact of the referenda can be isolated by fitting a Generalized Additive Model (GAM) to predict the daily spending and impressions. This model includes smooth functions for days until the referendum, the day of the year, and the trend, as well as a factor for the day of the week. Figure 5 shows the partial effects of the proximity of the referendum on the log-transformed expenditure and impressions. The results are consistent across all language groups and suggest a significant increase in ad spending and impressions 30 days before the vote. The 95% confidence intervals remain entirely above the baseline during this pre-referendum period, and then collapse immediately after the vote at $t=0$. These results confirm that ads are used intensively during the referendum campaign period.
@@
Figure 5: Partial effect of days to referendum (t=0) on log-transformed Expenditure (blue) and Impressions (orange), shown by language and aggregate. Shaded areas represent 95% confidence intervals.
@@
These findings lead to the next research question (RQ2.b): Given the increase in ad activity before referenda, is it possible to explore and measure the influence of referendum ads on the outcome of the vote? Figure 6 shows the distribution of all ad impressions supporting or opposing the referenda in relation to their outcomes. This figure suggests that there is a relationship between the number of impressions per side and referendum outcomes. When a referendum is rejected, the number of views of ads opposing the referendum looks to be higher. In contrast, if the outcome of a referendum is positive, the number of views of ads supporting it tends to be higher. Correlation analysis shows a moderate positive association between the ratio of Yes to No ad impressions and referendum approval outcome (Spearman $\rho=0.39$, $p=0.043$). This indicates that referendums with relatively more Yes-side advertising tended to be approved more often. Then, a Mann-Whitney U non-parametric test is used to independently compare the “Yes” and “No” impression volumes. The results suggest that approved referenda have a significantly higher volume of “Yes” impressions compared to rejected referenda ($U=146.0$, $p=0.0135$). In contrast, the volume of “No” impressions is not a statistically significant predictor of a referendum rejection ($U=117.0$, $p=0.1906$). In other words, greater exposure to pro-Yes advertising is associated with a higher probability of referendum approval, whereas exposure to pro-No advertising does not show a systematic link to rejection outcomes. It is important to note that this is a correlation and does not prove causation. In addition, popularity can act as a confounding factor: popular referendums are more likely to attract higher ad spending for the “Yes” side and are also more likely to pass. This is because “Yes” campaigns inherently face an uphill battle against the well-documented “status quo bias”, in which uncertain voters tend to vote “No” to avoid uncertainty [Bowler1998]. A “Yes” campaign must first have strong latent public support to overcome this bias and succeed. Thus, ad volume may indicate a well-funded campaign with this support rather than being an independent driver of the outcome [Gerber1998].
@@
4.3 Political issues
The third research question focuses on the federal elections in October 2023 and the issues promoted in ad campaigns. As described in 3.2.2, each ad published in the month before the federal elections is labelled with a relevance flag and with up to 3 keywords related to the topics discussed in the ad. Thanks to this data-processing procedure, it is possible to investigate whether parties discuss different issues during federal elections or promote different views on the same issues (RQ3.a).
@@ 4.3 Political issues
To further investigate similarities and differences between parties, a machine learning model is built to try to predict the party author of an ad (RQ3.b). For this multi-class classification task, a multinomial logistic regression model is trained on the ads relevant to the 2023 federal elections. Ads are weighted according to their total number of impressions to ensure that ads with a larger reach have a proportionally larger impact on the model’s parameters. For completeness, three different feature combinations are tested: (1) Demographic, representing the percentage of impressions per age and gender bucket; (2) Topic, which are one-hot encoded topic categories; and (3) a combination of Demographic + Topics.
@@
Demographic Topics Demographic
+Topics
@@
The recall-normalized confusion matrix (Figure 1(a)) highlights a more complex situation than the model’s high discriminative power might suggest. While recall is high for parties at the poles of the political spectrum, such as the SVP (0.78), it is notably low for the centrist parties Die Mitte and the GLP. This is particularly evident for Die Mitte, which achieves a recall of only 0.08. The confusion matrix shows that Die Mitte ads are most frequently misclassified as SVP (32%) and FDP (24%).
Figure 8: Marginal effects of the model with the best feature combination (demographic + topics). Only coefficients with a p-value below 0.05 are reported.
@@
5 CONCLUSIONS
Switzerland is renowned for its extensive use of direct democracy, which is characterized by frequent referendums and popular votes. Against this backdrop, this paper presents a large-scale, quantitative analysis of more than 40,000 political ads published on Facebook and Instagram in Switzerland. The dataset collected for this study comprises nearly 4.5 million Swiss francs in expenditures and more than 560 million impressions. This is a surprisingly high volume of activity given the relatively small Swiss electorate of 5.6 million.
The findings confirm that major political parties rely heavily on online advertising, using it not only to secure votes in federal elections, but also to influence referendum outcomes. The volume of ads is significantly higher before referenda and greater exposure to pro-Yes advertising has been shown to correlate statistically with referendum approval. In contrast, exposure to pro-No advertising does not show a systematic link to rejection outcomes.
Furthermore, there is significant evidence that political parties widely employ microtargeting based on demographics, such as age and gender. Generally, centrist and right-wing parties tend to attract older male audiences, whereas left-wing parties tend to focus on young women. A more granular analysis reveals that political parties adapt their microtargeting strategies to align also with Switzerland’s distinct linguistic regions. In fact, Switzerland is characterized by deep-seated heterogeneity, not only linguistically but also socio-economically, with significant disparities in wealth, industrial focus, and urban-rural landscapes. As literature has shown, Swiss heterogeneity is a primary driver of political cleavages [OeschRennwald] and voting behavior [SteenbergenMarco, CamatarriFavero], strongly correlating with economic inequalities and urbanity [Mantegazzi03072021]. Microtargeting analysis has several limitations imposed by the Meta Ad Library API, which provides an incomplete picture of targeting mechanisms. First, aside from the opacity of ad selection, the API only reveals targeting parameters related to geographical location, age, and gender. This omits the vast array of other attributes that Meta ad algorithm is known to use for profiling, such as international migration [ZagheniGummadi] or health-related behaviors [MejovaWeberBenevenuto]. Second, the available data does not distinguish between targeting decisions made manually by the campaign advertiser and those driven by Meta’s ad-delivery algorithms.
The content analysis of the ads reveals a clear pattern of issue divergence. Rather than engaging in direct debate with different views on the same issues, parties tend to focus their campaigns on distinct sets of topics, effectively “talking past each other”. This behavior aligns with the well-established political science theory of “issue ownership” [Petrocik], which posits that parties strategically emphasize topics where they have a long-standing reputation for competence [BudgeFarlie1983]. By selectively focusing on “their” issues, parties avoid competing on unfavorable ground and instead attempt to make their own issue priorities the most important for voters [amoros2013issue, BELANGER2008477]. Finally, the results confirm that the audience reach and topics promoted in advertisements differ significantly enough across parties to be used as effective features for a machine learning model to predict an ad’s author.
This study contributes to the ongoing debate about the impact of online political advertising on democratic processes. Although the potential consequences of this phenomenon are still being debated, censoring ads on social media platforms may not be preferable. Instead, transparency of data and algorithms, and data-driven analyzes, would empower users by providing them with the necessary information to understand how they are being targeted.
APPENDIX
Appendix A DATA
Table A1: Thematic Keyword Groups for Data Collection

### v2

@@ 3.2 Referenda
Figure 4 shows the seven-day moving average of total political ad views. The gray vertical lines represent referendums, and the inset shows the total number of views per language in the 30 days before the federal elections of October 22, 2023. Clearly, this event caused the most significant increase in views. At first glance, it is instead difficult to answer RQ2.a and determine whether advertising activity increases before the 14 referendum dates. As preliminary step, a Wilcoxon signed-rank test is applied to compare ad views and spending in the 30 days before and after each referendum. For completeness, this analysis is conducted separately for German, French, and Italian ads. The results of all the tests (shown in S4 Table) confirm a significant increase in ad views and spending in the 30 days before a referendum ($p<0.001$ in all cases).
@@
Although the Wilcoxon signed-rank test confirms a difference, it does not account for confounding factors, such as seasonality or other temporal patterns. The true impact of the referenda can be isolated by fitting a GAM to predict the daily spending and impressions. Figure 5 shows the partial effects of the proximity of the referendum on the log-transformed expenditure and impressions. The results are consistent across all language groups and suggest a significant increase in ad spending and impressions 30 days before the vote. The 95% confidence intervals remain entirely above the baseline during this pre-referendum period, and then collapse immediately after the vote at $t=0$. These results confirm that ads are used intensively during the referendum campaign period.
@@
Figure 5: Temporal dynamics of advertising intensity. Partial effects of the “days to referendum” variable on log-transformed Expenditure (blue curve) and Impressions (orange curve). The vertical dotted line at $t=0$ marks the day of the vote. The analysis isolates the referendum effect from seasonal trends and is presented for each linguistic region and the national aggregate. Shaded areas represent 95% confidence intervals.
@@
These findings lead to the next research question (RQ2.b): Given the increase in ad activity before referenda, is it possible to explore and measure the influence of referendum ads on the outcome of the vote? Figure 6 shows the distribution of all ad impressions supporting or opposing the referenda in relation to their outcomes. This figure suggests that there is a relationship between the number of impressions per side and referendum outcomes. When a referendum is rejected, the number of views of ads opposing the referendum looks to be higher. In contrast, if the outcome of a referendum is positive, the number of views of ads supporting it tends to be higher. Correlation analysis shows a moderate positive association between the ratio of Yes to No ad impressions and referendum approval outcome (Spearman $\rho=0.39$, $p=0.043$). This indicates that referendums with relatively more Yes-side advertising tended to be approved more often. Then, a Mann-Whitney U non-parametric test is used to independently compare the “Yes” and “No” impression volumes. The results suggest that approved referenda have a significantly higher volume of “Yes” impressions compared to rejected referenda ($U=146.0$, $p=0.0135$). In contrast, the volume of “No” impressions is not a statistically significant predictor of a referendum rejection ($U=117.0$, $p=0.1906$). In other words, greater exposure to pro-Yes advertising is associated with a higher probability of referendum approval, whereas exposure to pro-No advertising does not show a systematic link to rejection outcomes. It is important to note that this is a correlation and does not prove causation. In addition, popularity can act as a confounding factor: popular referendums are more likely to attract higher ad spending for the “Yes” side and are also more likely to pass. This is because “Yes” campaigns inherently face an uphill battle against the well-documented “status quo bias”, in which uncertain voters tend to vote “No” to avoid uncertainty [11]. A “Yes” campaign must first have strong latent public support to overcome this bias and succeed. Thus, ad volume may indicate a well-funded campaign with this support rather than being an independent driver of the outcome [33].
@@
3.3 Federal elections
The third research question focuses on the federal elections in October 2023 and the issues promoted in ad campaigns. As described in the Federal Election Subsection, each ad published in the month before the federal elections is labeled with a relevance flag and with up to 3 keywords related to the topics discussed in the ad. Thanks to this data-processing procedure, it is possible to investigate whether parties discuss different issues during federal elections or promote different views on the same issues (RQ3.a).
@@ 3.3 Federal elections
To further investigate similarities and differences between parties, a machine learning model is built to try to predict the party author of an ad (RQ3.b). For this multi-class classification task, a multinomial logistic regression model is trained on the ads relevant to the 2023 federal elections. For completeness, three different feature combinations are tested: (1) Demographic, representing the percentage of impressions per age and gender bucket; (2) Topic, which are one-hot encoded topic categories; and (3) a combination of Demographic + Topics.
@@
Demographic Topics Demographic + Topics
@@
The recall-normalized confusion matrix (Figure 8.a) highlights a more complex situation than the model’s high discriminative power might suggest.
Figure 8: Recall-normalized (a) and Precision-normalized (b) confusion matrices. The metrics refer to the Demographic + Topics model. The diagonal shows the per-class recall (a) or precision (b).
While recall is high for parties at the poles of the political spectrum, such as the SVP (0.78), it is notably low for the centrist parties Die Mitte and the GLP. This is particularly evident for Die Mitte, which achieves a recall of only 0.08. The confusion matrix shows that Die Mitte ads are most frequently misclassified as SVP (32%) and FDP (24%).
Figure 9: Marginal effects of the model with the best feature combination (demographic + topics). Positive values indicate a higher likelihood of an ad belonging to a specific party when that topic is present. Only statistically significant coefficients ($p<0.05$) are displayed.
@@
4 Discussion
The 2024 European Parliament Youth Survey [24] reports that 42% of European youth (ages 16–30) rely primarily on social media platforms such as TikTok, Instagram, and YouTube for political and social news. Consequently, the paid advertising and algorithmic mechanisms that govern visibility on these platforms now have a disproportionate influence on the democratic process. Switzerland’s system of direct democracy, characterized by frequent popular referendums, is particularly vulnerable to these shifts in democratic deliberation. It is little wonder that online political advertising has shifted from a peripheral tool to a central mechanism for influencing both elections and policy outcomes.
Against this backdrop, this article provides a comprehensive analysis of 40,000 political advertisements (560 million impressions) published on Meta from 2021 to 2025. A critical finding is the strong correlation between the volume of “pro-Yes” advertising and referendum approval. Unlike elections, where voters rely on party heuristics, referendums often involve complex technical issues in which the availability of information plays a decisive role. The correlation between higher ad exposure and favorable outcomes raises normative concerns. If visibility, purchased through an algorithmic auction, can influence policy decisions, then the “marketplace of ideas” could become a “marketplace of budgets”. Although our data reveals a correlation rather than a direct causal relationship, it suggests that well-funded interest groups may use algorithmic reach to circumvent traditional deliberation, especially in “low-salience” referendums where voters have fewer preconceived notions. Furthermore, evidence of demographic microtargeting confirms that political parties are actively segmenting the electorate, rather than merely broadcasting messages. Their adaptation of strategies to Switzerland’s linguistic and regional divisions demonstrates the sophistication of these campaigns. As previous studies have shown, Swiss heterogeneity is a primary driver of political divisions [48, 42] and voting behavior [57, 13]. Our results suggest that algorithms are effectively automating the exploitation of these societal fault lines. By delivering different narratives to young urban women than to older rural men, microtargeting reinforces existing socioeconomic and cultural divisions, which could lead to the polarization seen in other multicultural democracies.
The content analysis reveals a distinct pattern of divergent issues, where political actors “talk past each other”. Rather than engaging in a shared debate, parties retreat into safe topics, effectively creating parallel public spheres. This behavior aligns with the theory of issue ownership [51], in which parties emphasize topics on which they have a reputational advantage [12]. By selectively focusing on “their” issues, parties avoid competing on unfavorable ground and instead attempt to make their own issue priorities the most important for voters [2, 7]. Social media algorithms likely amplify this effect beyond what is possible in traditional media. In a television broadcast, viewers are exposed to the opposition’s arguments, even if they disagree with them. In the algorithmic environment observed, parties can target their “owned” issues exclusively to sympathetic demographics, insulating their base from counterarguments. This suggests that digital microtargeting transforms issue ownership from a campaign tactic into a mechanism of epistemic closure.
These conclusions should be interpreted in light of the limitations imposed by the Meta Ad Library API. The current transparency tools only reveal basic targeting parameters, such as age, gender, and location, and omit the vast array of behavioral and interest-based attributes used for profiling [62, 4]. Additionally, the data cannot distinguish between targeting decisions made by campaign managers and those optimized automatically by Meta’s delivery algorithms. This “black box” is a significant barrier to accountability. Without access to the precise logic of audience selection, researchers cannot determine whether elections are being manipulated by malicious actors or merely optimized by profit-driven algorithms.
This study contributes to the ongoing debate on the governance of the digital public sphere by providing a novel, data-driven analysis of the impact of political advertising on social media on direct democracy. Although censoring political ads entirely may be undesirable, maintaining the current level of secrecy is not an option. Empowering users and researchers requires shifting from voluntary transparency to comprehensive data access. Citizens can only move from being passive recipients of tailored messaging to active participants in a shared democratic process by understanding how they are being targeted.
5 Materials and methods
This study relies on data from two main sources: the Swiss Federal Statistical Office [^1] https://abstimmungen.admin.ch/en/overview and the Meta Ad Library API [^2] https://www.facebook.com/ads/library/api/. The Federal Statistical Office provides detailed statistics related to referenda, including results by canton, participation percentages, and full descriptions of referendum questions.
5.1 Meta Ad Library
The Meta Ad Library API discloses advertisements about social issues, election, or politics that were delivered on Facebook or Instagram during the past seven years. Although the API allows for specifying search parameters, such as limiting results to a specific country, all queries must contain keywords or author names. The following information can be extracted for each advertisement: the publication date; the end time of the campaign; the textual content; and the number of impressions (provided as an interval). The term “impressions” refers to non-unique visualizations. It is also possible to access the campaign expenditure (again presented as an interval) and the audience’s demographic profile, specifically their gender (male, female, or unknown) and age (categorized into seven buckets). In line with existing literature [16], the midpoint of the interval’s boundaries is computed for values presented as ranges (expenditure and impressions). For unbounded ranges, only the defined boundary is used. To quantify the extent of demographic microtargeting, two statistical measures are employed. First, gender disparities are analyzed using the odds ratio (OR). This metric measures the likelihood of an advertisement being shown to a male user versus a female user. An OR of $1.0$ indicates a perfectly balanced gender distribution; values greater than $1.0$ indicate a bias toward a male audience, while values less than $1.0$ indicate a bias toward a female audience. Second, differences in age targeting are evaluated using the Wasserstein distance (WD), often referred to as the Earth Mover’s Distance. Unlike comparing simple averages, the WD accounts for the entire shape of the probability distributions. It quantifies the minimum amount of “work” required to transform one distribution into another. In the context of this study, the WD effectively represents the mean absolute difference in years between the age profile of a specific party’s audience and that of the comparison group (e.g., the rest of the dataset).
The collection is initiated using a broad set of election-related keywords (see Table 2) in German, French, and Italian due to the API’s requirement for keyword- or author-based searches. This process yielded an initial corpus of 40,044 ads (22,494 from German keywords, 10,794 from French, and 6,756 from Italian) published between January 1, 2021, and September 30, 2025. The API’s country filter includes advertisements displayed in Switzerland, even if they are intended for a different audience. To create a Swiss-centric dataset, ads are further filtered to include only those for which more than 50% of impressions come from users based in Switzerland. This reduces the dataset to 34,559 ads from 2,246 unique pages, which received an estimated 560 million impressions and cost over CHF 4.5 million.
Although the API provides a page name (username) for each ad, this identifier is frequently ambiguous or does not directly correspond to a particular political party, even when the content of the page clearly promotes a party’s agenda. As shown in Figure 4, the 2023 federal elections represent the most significant activity peak in the dataset. For this reason, the 1,224 pages that published at least one ad in the 30 days preceding the federal election are manually labeled to establish whether they are related to a specific party. This process identified 943 pages that could be directly attributed to a political party, while the remaining 281 pages belong to entities such as media organizations (newspapers and daily newspapers), NGOs, neighborhood committees, and non-partisan referendum groups. Of the 943 political pages, 17 major parties and smaller entities (e.g., municipal parties and civic lists) are identified. However, only six parties — SVP, SP, FDP, Die Mitte, GRÜNE, and GLP — play a dominant role in the advertising landscape, accounting for 95% of all impressions. Collectively, these parties also secured 91.84% of the votes in the 2023 federal elections. Tables 34, and 5 provide detailed statistics on impressions, volume, and ad expenditure broken down by party and language.
5.2 Content Annotation
5.2.1 Referenda
To evaluate the intensity of advertising activity, a Wilcoxon signed-rank test is employed to determine if there is a significant increase in the number of ads and expenditure in the period preceding a referendum. The Wilcoxon signed-rank test is a non-parametric statistical hypothesis test used to compare two related samples, matched samples, or repeated measurements on a single sample to assess whether their population mean ranks differ. In this study, it is used to compare the 30-day “pre-referendum” window against the 30-day “post-referendum” window for each vote. This test is particularly suitable because it does not assume a normal distribution of the data, which is necessary given the highly skewed nature of advertising impressions and expenditure.
However, the Wilcoxon test only identifies a difference between two time blocks and does not account for external confounding factors such as seasonality, day-of-the-week effects, or long-term growth trends in the dataset. To isolate the specific impact of referendum proximity, a Generalized Additive Model (GAM) is deployed. A GAM is a flexible regression technique that extends Generalized Linear Models by allowing the linear predictor to depend on unknown smooth functions of some predictor variables. By using smoothers for the “days until referendum”, the day of the year, and the general trend, the model can separate the temporary spike caused by a political campaign from recurring temporal patterns.
5.2.2 Federal Election
The second analysis focuses on the 2023 federal elections. Unlike a referendum, which addresses one or a few political issues, federal elections involve political parties promoting broader agendas. Therefore, it is necessary to not only identify election-relevant ads but also determine the specific topic(s) they cover. Consistent with the previous task, only ads with at least 10 words published within 30 days before the federal election are selected. The model checkpoint and the API query date are the same of the previous task. The model’s output includes a Boolean parameter identifying whether the ad is relevant to the federal election, as well as up to three English keywords related to the ad’s topic. This initial extraction process yielded a large set of 739 unique keywords, which require consolidation into fewer categories. To this end, each keyword is encoded into a semantic vector using OpenAI’s state-of-the-art text-embedding-3-large model. These vectors are then clustered into 25 topics using the K-Means clustering algorithm [44]. This unsupervised method partitions the semantic vectors into $K=25$ distinct clusters by iteratively assigning each data point to the nearest centroid, thereby minimizing the within-cluster variance. To ensure reliability, qualitative validation is performed on a random subset of 100 ads. In this review, human annotation agreed with model-generated keywords in 96% of the cases.
To address RQ3.b, a multinomial logistic regression model is employed to predict the political party responsible for a given ad. This method generalizes binary logistic regression to multi-class classification problems by using the softmax function to model the probability distribution over $K$ possible outcomes (parties). For a given input vector $\mathbf{x}$ (containing demographic or topic features), the probability that the ad belongs to party $k$ is given by:
$P(y=k\mid\mathbf{x})=\frac{e^{\mathbf{w}_{k}^{T}\mathbf{x}}}{\sum_{j=1}^{K}e^{\mathbf{w}_{j}^{T}\mathbf{x}}}$ (1)
where $\mathbf{w}_{k}$ is the weight vector associated with class $k$. Ads are weighted according to their total number of impressions to ensure that ads with a larger reach have a proportionally larger impact on the model’s parameters.
A multinomial logistic regression model provides interpretable coefficients that reveal the relationship between specific features (e.g., a specific topic) and the likelihood of an ad originating from a specific party. An ablation study is conducted to evaluate feature importance using three configurations: only demographic features, only topic features, and a combination of the two.
The model’s performance is evaluated using the Area Under the Receiver Operating Characteristic curve (AUC ROC). The ROC curve plots the true positive rate (sensitivity) against the false positive rate (1-specificity) at various threshold settings. The AUC represents the probability that the model ranks a randomly chosen positive instance higher than a randomly chosen negative instance. Since this is a multi-class problem, the AUC is calculated using a One-vs-Rest strategy, where the performance for each party is evaluated against all other parties combined. An AUC score of 0.5 corresponds to random guessing, indicating a lack of discriminative ability. In contrast, a score of 1.0 signifies perfect classification. To better understand the model’s specific strengths and weaknesses, the recall and precision confusion matrices are also analyzed. Recall quantifies the model’s ability to correctly identify all ads belonging to a specific party (True Positive Rate). Precision, on the other hand, measures the accuracy of the model’s positive predictions for a given party (Positive Predictive Value).
A multinomial logistic regression model provides interpretable coefficients that reveal the relationship between specific features (e.g., a specific topic) and the likelihood of an ad originating from a specific party. An ablation study is conducted to evaluate feature importance using three configurations: only demographic features, only topic features, and a combination of the two.
The model’s performance is evaluated using the Area Under the Receiver Operating Characteristic curve (AUC ROC). The ROC curve plots the true positive rate (sensitivity) against the false positive rate (1-specificity) at various threshold settings. The AUC represents the probability that the model ranks a randomly chosen positive instance higher than a randomly chosen negative instance. Since this is a multi-class problem, the AUC is calculated using a One-vs-Rest strategy, where the performance for each party is evaluated against all other parties combined. An AUC score of 0.5 corresponds to random guessing, indicating a lack of discriminative ability. In contrast, a score of 1.0 signifies perfect classification. To better understand the model’s specific strengths and weaknesses, the recall and precision confusion matrices are also analyzed. Recall quantifies the model’s ability to correctly identify all ads belonging to a specific party (True Positive Rate). Precision, on the other hand, measures the accuracy of the model’s positive predictions for a given party (Positive Predictive Value).
6 Acknowledgments
I thank Dr. Dirk Helbing for fruitful discussions.
Data Availability
Data and code are available in the Zenodo repository at https://doi.org/10.5281/zenodo.18256330.
7 Supplementary Material
Table 2: Thematic Keyword Groups for Data Collection